# Neural Geometry Processing via Spherical Neural Surfaces

### v1

@@ 5. Spectral Analysis using SNS
Figure 10. Spectral SNS Modes across Isometric Poses. If we compute the first few eigenfunctions on SNSs that represent near-isometric surfaces, and order the functions according to their Rayleigh quotients, we see a clear correspondence between corresponding eigenfunctions on each surface. Although our computation of the LBO and its spectrum involves extrinsic terms such as mean curvature and normals, the Laplace Beltrami operator is intrinsic so its spectrum does not depend on the particular embedding of a surface into $\mathbb{R}^{3}$, therefore the eigenmodes of isometric SNSs should appear similar (up to a possible sign change).
Figure 11. Robustness of Neural LBO. We demonstrate that our optimized eigenfunctions are largely unaffected by differences in mesh density on the starting mesh, because we work directly with a smooth surface. Here we show that while the cotan discretization can be very inconsistent across changes in mesh density, the neural LBO is more consistent when applied to a scalar field defined on an SNS that is generated from different meshes of the same surface.
6. Evaluation
Implementation Details
@@ 6. Evaluation
All of our networks in our experiments were trained on Nvidia GeForce P8 GPUs. The SNS network is a Residual MLP with input and output of size three, and eight hidden layers with 256 nodes each. The networks used to represent scalar fields (in the eigenfunction and heat flow experiments) are very small Residual MLPs with input size three, output size three and two hidden layers with 10 nodes each. (To make the output value a scalar, we take the mean of the three output values.) For both scalar fields and SNS, the activation function is ‘Softplus’. We use the RMSProp optimizer with a learning rate of $10^{-4}$ and a momentum of 0.9. We trained each SNS for up to 20,000 epochs (eight-to-ten hours), and we fixed the normals regularization coefficient to be $10^{-4}$. The time per epoch increases slightly with the number of vertices in the mesh, but for simpler shapes (such as the sphere) the optimization converges in fewer epochs.
Eigenfunction Optimization
@@
Comparison against Analytic Functions
For some of our SNS computations, there exist special cases for which an exact analytic solution can be calculated. To test our estimates of curvature and principal curvature directions, we overfitted an SNS to a mesh of the closed surface with the parametrization
(28) $\mathbf{r}(\theta,\phi)=(1+0.4\sin^{2}\theta\sin(6\theta)\sin(6\phi))\begin{pmatrix}\sin\theta\cos\phi;\sin\theta\sin\phi;\cos\theta\end{pmatrix}.$
We used the Matlab Symbolic Math Toolbox™ to compute the normals, mean curvature, Gauss curvature and principal curvature directions in closed form. In Figure 3, we show these differential quantities on the mesh of the analytic surface and on the SNS. The computed curvatures align extremely closely, therefore the overfitting quality is very high and the curvature computations are accurate. There is a slight discrepancy at the poles, because we have used a global analytic parametrization which not valid where $\theta\in\{0,\pi\}$. On the other hand, in the SNS formulation we use selective co-ordinate relabelling to avoid issues related to using spherical polar co-ordinates close to the poles.
For evaluation of our optimized eigenfunctions, we overfitted an SNS to a sphere-mesh with 10,242 vertices and optimized for the first five non-constant eigenfunctions. The eigenspaces of the LBO on a sphere are the spaces spanned by the spherical harmonics of each frequency (refer to (Jarosz, 2008)). The first three spherical harmonics have eigenvalue -2 and the following five spherical harmonics have eigenvalue -6. Although the basis for each eigenspace is not unique, the first three eigenfunctions should also have an eigenvalue of -2 and the following five eigenfunctions should have an eigenvalue of -6, meaning that the Rayleigh Quotients should be equal to 2 and 6 (see Figure 4). We computed the ground-truth eigenvalues by manually integrating the Cartesian formulas (Jarosz, 2008) and we estimated the integral of the neural eigenfunctions using 100k Monte Carlo samples. Estimates were within $1\%$ of the ground-truth, and four out of five results were within $0.5\%$.
Neural LBO on Scalar Fields
Finally, we computed the result of applying the cotan LBO to the scalar field (sampled at the vertices of the dense icosphere mesh). At this resolution, all three versions of the LBO are extremely close, however, we notice that the cotan estimate on the ‘spike ball’ (column 2) contains a small discontinuity, not present in the SNS estimates. From a smooth shape and a smooth scalar field, the resulting $\Delta_{S}$ should be smooth, which is the case for SNS estimates.
Heat Flow and Mean Curvature Flow
@@
Similarly, we approximate a Mean Curvature Flow of an SNS by updating the SNS to $S_{\theta}-dH\mathbf{n}$ at each time step, and finetuning for up to 100 epochs. The results in Figure 8-right show the spiky analytic shape becoming closer to a sphere, over 150 iterations.
Figure 12. Handling Neural Implicits. Spherical Neural Surfaces can be generated from other neural representations, such as DeepSDF. Here, we have produced an SNS from a DeepSDF representation for a camera, by overfitting the SNS to a low-resolution marching-cubes mesh reconstruction and then finetuning this SNS to better align with the SDF.
Neural Implicits
To demonstrate robustness to the input representation, we provide an illustratory example of an SNS generated from a DeepSDF (Park et al., 2019), and show the normals and principal curvature directions on its SNS. We took a pretrained DeepSDF network and the optimized latent code for the camera shape (Comi, 2023). Then, as initialization, we overfitted an SNS to a coarse marching-cubes mesh. Then we took a set of samples on the SNS and projected them to deepSDF representation defined by the signed distance function (SDF), by moving them in the direction of the gradient of the SDF (which approximates the surface normal) by the signed distance at that point, and use them to finetune the SNS.
7. Conclusion
We have presented spherical neural surfaces as a novel representation for genus-0 surfaces along with supporting operators in the form of differential geometry estimates (e.g., normals, First and Second Fundamental Forms), as well as surface gradients and surface divergence operators. We also presented how to compute a continuous Laplace Beltrami operator and its lowest spectral modes. We demonstrated that ours, by avoiding any unnecessary discretization, produces robust and consistent estimates even under different meshing concerning vertex sampling as well as their connectivity.
@@ 7. Conclusion
Since our SNSs rely on a sphere for the parametrization, it limits us to genus-0 surfaces. If we could choose from a range of canonical surfaces (e.g., torus), this would allow us to seamlessly process higher genus shapes and, possibly, to produce neural surfaces without such high distortions. Another option would be to rely on local parametrizations, but then we would have to consider blending across overlapping parametrizations.
@@
A severe limitation of our current realization is its long running time. While we expect that better and optimized implementations would increase efficiency significantly, we also need to make changes to our formulation. Specifically, our current spectral estimation is sequential, which is not only slow, but leads to an accumulation of errors for later spectral modes. Hence, in the future, we would like to jointly optimize for multiple spectral modes.
@@ Acknowledgements
This project has received funding from the UCL AI Centre. RW wishes to thank Johannes Wallner for his constructive comments regarding Section 4.

### v2

@@ 5 Spectral Analysis using SNS
Figure 9: Spectral SNS Modes across Isometric Poses. If we compute the first few eigenfunctions on SNSs that represent near-isometric surfaces, and order the functions according to their Rayleigh quotients, we see a clear correspondence between corresponding eigenfunctions on each surface. Although our computation of the LBO and its spectrum involves extrinsic terms such as mean curvature and normals, the Laplace Beltrami operator is intrinsic so its spectrum does not depend on the particular embedding of a surface into $\mathbb{R}^{3}$, therefore the eigenmodes of isometric SNSs should appear similar (up to a possible sign change).
Figure 10: Robustness of Neural LBO (Analytic). We demonstrate that our neural LBO is largely unaffected by differences in mesh density and mesh quality on the starting mesh. While the common mesh-based LBOs are sensitive to the quality and the density of the mesh, the neural LBO is more consistent when applied to a scalar field defined on an SNS that is generated from different meshes of the same surface. The mesh-based LBO is the cotan LBO for the triangle-meshes and the virtual-refinement Laplace [BHKB20] for polygonal meshes. The scalar field is analytic (a variable frequency sine-wave) on a sphere (for which the mean-curvature is one and the normals are trivial), because this allows us to compute hessian, gradient, normals and mean curvature analytically, so that we can compare against the ground truth LBO using equation 16. Please refer to the supplemental for further details. The error colourmap runs from white (no error) to red (higher error).
Figure 11: Robustness of Neural LBO (Non-Analytic). For non-analytic surfaces, we do not have access to a ground-truth LBO, because the precise underlying surface geometry is ambiguous. However, we show that the neural LBO is more consistent than mesh-based LBOs when fitted to different meshes of the same surface. This is because the smooth geometry of the SNS remains quite consistent when fitted to different meshes of the same surface, and the neural LBO is defined everywhere on the smooth surface - not just at the mesh-vertices. The neural LBO on mesh 2 does display some visible differences, however this is because the hair of Igea contains some high-frequency details. This is a signal which was lost in the lower-resolution meshes, so it is not an artefact or noise.
6 Evaluation
Dataset.
We evaluate our method on various meshes of the unit sphere, various meshes of the Igea model, and four analytic shapes. In order to evaluate the robustness of our method, we included several ‘bad’ meshes (e.g. with many thin triangles) and unusual meshes such as those with a very sharp contrast in triangle size across different regions, and two polygonal meshes. (See Figures 10and 11.)
The four analytic shapes, which we designed for this experiment, are: mushroom, flower, bobble, and star. In Figure 6, show results on the mushroom only. Refer to the supplemental, for the analytic parametrizations of all four shapes, and for additional results on the flower, bobble, and star. The advantage of using analytic shapes for evaluation is that we can compute the ground truth for any derived geometric quantity (such as curvature). We used symbolic Matlab to compute the ground truth differential quantities, given the analytic parametrization. We provide text files, in the supplemental, which contain the symbolic expressions.
Evaluation Metrics.
To evaluate the quality of the surface representation, we measure:
  1. (i)
    The distance from each point on the reconstructed surface, to the analytic surface (geometric error).
  2. (ii)
    The angle deviation of the surface normal at each (reconstructed) surface point, compared to the ground truth surface normal at the the closest analytic surface point.
  3. (iii)
    The angle deviation of the minimum curvature direction at each (reconstructed) surface point, compared to the ground truth minimum curvature direction at the the closest analytic surface point.
  4. (iv)
    The absolute error of the Gaussian curvature and Mean curvature at each (reconstructed) surface point, compared to the ground truth curvatures at the the closest analytic surface point.
For the Laplace Beltrami operator, we compare our LBO to other LBOs when applied to a frequency-varying sine-wave type scalar field, and in the case of a sphere, we compute the absolute error with respect to the analytic LBO (see supplemental for details).
Comparison Methods.
We compare the SNS against three alternative methods for estimating geometric quantities. (These methods do not support LBO computation or spectral analysis.)
  • Monge-fitting [PC24]] is a classical mesh-based method that uses osculating jets to approximate differential quantities of the underlying curved surface. We used the official implementation available via CGAL.
  • NFGP [YBHK21] is a method has similar motivation to ours, in the sense that they fit a neural representation to a mesh and compute differential quantities in order to perform edits in the shape. However, they use a deep-implicit representation, so their differential quantities additionally depend on the representation’s adherance to the Eikonal property, whereas we use an explicit representation.
  • i3D [NSS22] is a deep implicit method also computes differential quantities, and improves the overfitting process itself by encouraging the curvature of the deep implicit to match the computed curvature of the input mesh. Because our method, NFGP, and Monge fitting do not require the curvature information from the mesh, we did not provide the mesh-curvature information to this method either and we found that the method suffered as a result.
(i) What is the effect on the estimates when we fit the SNS to different resolution meshes of an analytic shape?
In Figure 6, we fit SNSs to analytic shapes (mushroom), and plot the absolute error of derived geometric quantities (against the ground truth, which is computed from the analytic parametrization using symbolic Matlab). We fit SNSs to a coarse, medium and fine mesh (with 2562, 10242 and 40962 vertices, respectively) and, for reference, we also fit an SNS directly to the analytic parametrization (without meshing).
As expected, the accuracy of the derived quantities is highest when we fit the SNS directly to the analytic parametrization, and the accuracy of the mesh-fitted SNSs approaches the analytically-fitted SNS as the mesh resolution increases. Since the accuracy of the SNS fitted to the fine-resolution mesh (40962 vertices is consistently close to the accuracy of the analytically-fitted SNS, we argue that fitting the SNS to a reasonably fine mesh is a good way to find a smooth approximation of an underlying smooth surface, when no analytic parametrization is available. We also show the distribution of errors for one of the shapes (mushroom), for SNSs fitted to each resolution of the mesh, and for the three comparison methods: Monge-fitting, NFGP and i3D.
(ii) How do the position, normals and curvature estimates of SNS compare to those from other methods?
In Figure 6, alongside the SNS errors, we show the absolute errors of computing the same geometric quantities on the fine-resolution mesh, using three alternative methods. The first method is Monge-fitting, which fits a piecewise polynomial surface to the mesh and allows us to compute the geometric quantities on the vertices. The other two methods (i3d and NFGP) are deep implicit methods, which fit a neural signed-distance function (a deepSDF) to the mesh. The i3D method has the option to use mesh-curvature information to improve the fitting, but for consistency, we did not provide this, because the other neural methods (ours, and NFGP) do not require additional mesh-curvature information. We use the CGAL implementation for Monge-fitting, and the original author -implementations for the two deep-implicit methods. The geometric quantities on a deepSDF are computed using automatic differentiation. (Although this appears at first glance to be similar to computing geometric quantities on an SNS, the computations are different due to deepSDF being an implicit representation while SNS is explicit. In contrast to SNS, deepSDFs use a soft constraint enforced by an Eikonal loss term, without which the deepSDF can only be used for mesh-reconstruction and not for the estimation of geometric quantities.)
We observe that given only a fine-resolution mesh, the most accurate method is consistently the SNS, except on the flower shape where it is Monge-fitting. The second most accurate method appears, on average, to be the Monge-fitting, but impressively the medium-resolution SNS is sometimes able to perform similarly well, showing that SNS has can robustly interpolate from limited geometric data.
On all shapes, i3D performed worse than the fine-resolution SNS and the Monge-fitting. The NFGP method produces consistently poor results. We believe that this method struggles to capture finer surface details (e.g. on the mushroom) and thin structures (e.g. on the flower), at least with the default settings. However, we were limited by compute resources and it is possible that better results could be achieved given a longer training time (please see the ‘NFGP Note’ in the Supplemental). Observations of the reconstructed meshes show that neither of the deep implicit methods capture the input geometry well, leading to poor results in the estimation of geometric quantities.
(iii) How consistent is the SNS LBO, when we fit an SNS to different meshes of the same shape?
In Figure 10 and 11, we put a test-function (a variable-frequency sine-wave scalar field) on two different surfaces: sphere and Igea, respectively. We fit the SNS to various different meshings of these shapes, and in each case compute the result of the applying the SNS LBO to the scalar field. To fit an SNS to a polygonal mesh, we first convert to a triangle mesh by splitting the faces.
For reference, we also show the result of applying a mesh-based LBO to the scalar field (sampled at the vertices). For triangle meshes, we use the cotan LBO (with vertex area as $1/3$ the one-ring area, and a lumped mass matrix). For polygonal meshes, we use the Virtual Refinement Laplace operator from Polygon Laplacian Made Simple [BHKB20], with the default Area Minimizer setting.
Visually, the SNS LBO produces more consistent and less noisy results across different meshings. This is in part because the SNS LBO is discretization-free, so we can sample at any point (not just vertices of the original mesh), and partly because the SNS LBO is able to leverage information from the derivatives of the scalar field (which is always available when the scalar field is analytic, or neural) whereas the mesh-based methods cannot use this additional available information.
(iv) How does the SNS LBO compare to the ground truth LBO, when the analytic operator is available?
In the case of a sphere, we know the mean curvature and normals analytically at every point, so we can use Equation 16 to compute the ground-truth result of applying the LBO to the (analytically-defined) scalar field. We show this result in Figure 10 , and also display the error maps of the mesh-based LBOs and the neural LBO. Clearly, the error of the neural LBO is smaller than the error of the mesh-based LBO, and the error is smallest where the scalar field is low-frequency.
(v) Neural LBO on Scalar Fields on non-analytic Surfaces.
Finally, we computed the result of applying the cotan LBO to the scalar field (sampled at the vertices of the dense icosphere mesh). At this resolution, all three versions of the LBO are extremely close, however, we notice that the cotan estimate on the spike ball (column 2) contains a small discontinuity, not present in the SNS estimates. From a smooth shape and a smooth scalar field, the resulting $\Delta_{S}$ should be smooth, which is the case for SNS estimates.
Implementation Details.
@@ 6 Evaluation
All of our networks in our experiments were trained on Nvidia GeForce P8 GPUs. The SNS network is a Residual MLP with input and output of size three, and eight hidden layers with 256 nodes each. The networks used to represent scalar fields (in the eigenfunction and heat flow experiments) are very small Residual MLPs with input size three, output size three and two hidden layers with 10 nodes each. (To make the output value a scalar, we take the mean of the three output values.) For both scalar fields and SNS, we use softplus as the activation function. We use the RMSProp optimizer with a learning rate of $10^{-4}$ and a momentum of 0.9. We trained each SNS for up to 20,000 epochs (eight to ten hours), and fixed the normal regularization coefficient to be $10^{-4}$. The time per epoch increases slightly with the number of vertices in the mesh, but for simpler shapes (such as the sphere) the optimization converges in fewer epochs.
Can our optimization process reproduce eigenfunctions of the LBO when the analytic solution is known?
@@
Figure 12: Heat Flow (top) and Mean Curvature Flow (bottom). (Top) We can evolve a given scalar field, $f$, specified over the surface of an SNS, $S_{\theta}$, using the Partial Differential Equation (PDE) $f_{t}=\Delta_{S_{\theta}}f$ (the heat equation). We represent the evolving scalar field as a small MLP, whose weights are ‘finetuned’ for up to 100 epochs after every time step. Darker colors denote low values (cold) with lighter colors being high (hot). (Bottom) Taking advantage of the differentiable nature of our representation, we can also compute a mesh-free Mean Curvature Flow (MCF) in which the Mean Curvature and normals are updated at every iteration (the co-ordinate function evolves according to the PDE $X_{t}=\Delta_{S_{\theta(t)}}X$). We update the Spherical Neural Surface using up to 100 finetuning steps, after every iteration of the flow. This formulation of MCF prevents singularities from forming, without any special handling to prevent them [KSBC12].
Heat Flow and Mean Curvature Flow.
@@
Similarly, we approximate a Mean Curvature Flow of an SNS by updating the SNS to $S_{\theta}-dH\mathbf{n}$ at each time step, and finetuning for up to 100 epochs, so that the surface evolves according to the PDE $X_{t}=\Delta_{S_{\theta(t)}}X$ The results in Figure 12-right show the spiky analytic shape becoming closer to a sphere, over 150 iterations.
Figure 13: Handling Neural Implicits. Spherical Neural Surfaces can be generated from other neural representations, such as DeepSDF. Here, we have produced an SNS from a DeepSDF representation for a camera, by overfitting the SNS to a low-resolution marching-cubes mesh reconstruction and then finetuning this SNS to better align with the SDF.
Neural Implicits.
To demonstrate robustness to the input representation, we provide an illustratory example of an SNS generated from a DeepSDF [PFS19], and show the normals and principal curvature directions on its SNS. We took a pretrained DeepSDF network and the optimized latent code for the camera shape [Com23]. Then, as initialization, we overfitted an SNS to a coarse marching-cubes mesh. Then we took a set of samples on the SNS and projected them to deepSDF representation defined by the signed distance function (SDF), by moving them in the direction of the gradient of the SDF (which approximates the surface normal) by the signed distance at that point, and use them to finetune the SNS.
7 Conclusion
We have presented spherical neural surfaces as a novel representation for genus-0 surfaces. Our smooth and differentiable neural surface representation inherently supports the computation of continuous differential geometric quantities and operators such as normals, First and Second Fundamental Forms), surface gradients and surface divergence. We also presented how to compute a continuous Laplace Beltrami operator and its lowest spectral modes.
We demonstrate that SNSs, by avoiding the pitfalls of discretization, produce robust and consistent differential geometric estimates for surfaces both across meshing parameters such as sampling and vertex connectivity, and when compared to alternate neural and mesh-based surface representations.
@@ 7 Conclusion
Since our SNSs rely on a sphere for the parametrization, it limits us to genus-0 surfaces. We chose the sphere as a domain, because it allows us to process many common shapes without needing to ‘cut’ them into a disc-topology (as shown in, e.g., Neural Surface Maps [MAKM21]). If we could choose from a range of canonical surfaces (e.g., torus), this would allow us to seamlessly process higher genus shapes and, possibly, to also produce neural surfaces without such high distortions. Another option would be to rely on local parametrizations, but then we would have to consider blending across overlapping parametrizations.
Solving Variational Problems
We are optimistic that the techniques demonstrated in our work have the potential to allow us to solve other variational problems, beyond the example of computing a spectral basis via energy minimization. We would like to know whether similar energy minimizations may be used on neural surface-paths, to compute geodesics, and on scalar- and vector-fields, to compute heat-flow/mean-curvature flow without fine-tuning the network at every step.
@@
A limitation of our current realization is its long running time. While we expect that better and optimized implementations would increase efficiency significantly, we also need to make changes to our formulation. Specifically, our current spectral estimation is sequential, which is not only slow, but leads to an accumulation of errors for later spectral modes. Hence, in the future, we would like to jointly optimize for multiple spectral modes.
@@ Acknowledgements
The authors wish to thank Johannes Wallner, for his constructive comments on an early version of this work, as well as Uday Kusupati and Karran Pandey, for proofreading. We also thank Morena Protti, for the amazing treefrog mesh. RW was supported by the Engineering and Physical Sciences Research Council (grant number EP/S021566/1).